what the follwing system of equation by substitution and check. x-y=13 9x-y=45
Answer by mananth(16946)

9 x + -1 y = 45 .............1
-1 y = 45 + -9 x
/ -1 =
y = -45 + 9 x
1 x + -1 y = 13 .....................2
Substitute y in (2)
x-1(-45+9x)=13
x+45-9x=13
x-9x=-32
-8x=-32
/ -8
x=4
Plug the value of x in (1)
9x-y = 45
9 * 4 + -1 y = 45
36 + -1 y = 45
-1 y = 9
/ -1
y= -9
CHECK
x-y=13
4-(-9)=13

m.ananth@hotmail.ca